# A Unified Photometric Redshift Calibration for Weak Lensing Surveys using the Dark Energy Spectroscopic Instrument

### v1

@@ 6.1 Direct Calibration
6.2 KiDS-1000 Results
Whereas our results for DES Y3 and HSC Y1 are in good agreement with the fiducial results presented in Myles et al. (2021) and Hikage et al. (2019), we find stronger differences with the KiDS-1000 results presented in Hildebrandt et al. (2021), as indicated in Fig. 6 and Table 1.
One major difference in our analysis is the use of high-quality photometric redshifts from the COSMOS2020 catalog to supplement spectroscopic data. Using COSMOS2020 redshifts is necessary since some galaxies are severely under-represented in the DESI COSMOS-XMM catalog. The effect of including COSMOS2015 multi-band redshifts in the KiDS-1000 analysis was studied in van den Busch et al. (2022) using the same methodology as in Hildebrandt et al. (2021). It was found that a calibration that includes COSMOS2015 redshifts leads to higher mean redshifts even for identical SOM cells. Since we also find higher mean redshifts on average, this indicates that the use of COSMOS2020 redshifts may be one of the reasons for the different results we find here. In turn, part of this may be driven by scatter and catastrophic outliers in COSMOS2020 redshifts (Weaver et al., 2022).
As mentioned earlier, in principle, assigning extra weights to galaxies in the calibration field to reproduce the color and magnitude distribution in the entire survey is not necessary for unbiased results. This requires that the color distribution and depth of the calibration field is representative of the full survey. However, for KiDS-1000, we find that the KiDS observations in the COSMOS field are somewhat deeper than the wider survey. Whereas Hildebrandt et al. (2021) implicitly make the KiDS COSMOS catalog reproduce the color and magnitude distribution in KiDS-1000 via a SOM, we did not. This may also explain some of the differences in our results. However, it is not clear that applying extra weights, as was done in Hildebrandt et al. (2021), resolves all issues associated with non-representative photometry. In the appendix, we present a comparison of the color and magnitude distribution of galaxies in the KiDS COSMOS catalog compared to the wider survey.
Finally, another reason for the different results may be related to the SOM cell quality cuts that were applied in the KiDS-1000 analysis (Hildebrandt et al., 2021). Before performing the photometric redshift calibration, Hildebrandt et al. (2021) defined the so-called gold selection by removing galaxies for which the mean spectroscopic redshifts of the calibration sources differ systematically from the mean photometric redshifts in that SOM cell.
@@ 6.2 KiDS-1000 Results
In the above equation, the subscripts d and w denote averages of the deep calibration field and the wide survey, respectively. Also, whereas the quality cut is applied to each SOM cell individually, the normalized median absolute deviation on the right-hand side is taken over all SOM cells. The goal of this quality cut is to remove parts of the color and magnitude space for which photometric redshifts are less reliable. Using the MICE mock catalogs, Wright et al. (2020) show that this is indeed the case and reduces the bias of the SOM direct calibration method.
@@
These advantages notwithstanding, the quality cut may lead to slight statistical biases. Whether any SOM cell fulfills this quality cut is a random variable and can be very stochastic given that one typically has $\mathcal{O}(10)$ galaxies with spectroscopic redshifts in each SOM cell. A potential issue is that Hildebrandt et al. (2021) use the same galaxies to define the quality cut and to perform the photometric redshift calibration. In particular, by construction, all SOM cells in the gold selection pass the quality cut, i.e., do not show significant differences between the spectroscopic and photometric redshifts. For example, the analysis of Hildebrandt et al. (2021) finds that for the first tomographic bin, KiDS-1, no single SOM cell has a mean redshift above $\sim 0.6$ which implies that the reconstructed $n(z)$ drops off sharply for $z\gtrsim 0.6$. However, this is, to some extent, a statistical artifact from the fact that the same sample was used to determine the quality cuts and to perform the calibration. If one uses galaxies not used to define the quality cut, as was done in this work, it is possible that individual SOM cells may have mean redshifts above $\sim 0.6$. This will lead to a higher $n(z)$ at high redshifts, as we see in the comparison of the reconstructed $n(z)$ for KiDS-1. In summary, the implementation of the quality cuts in Hildebrandt et al. (2021) may introduce subtle statistical biases which may also be partially responsible for the different results.
@@
Figure 7: Posterior constraints on $S_{8}$ and $\Omega_{\mathrm{m}}$ from the KiDS-1000 cosmic shear analysis of COSEBIs. In blue, we show the fiducial results from Asgari et al. (2021), whereas we show the updated results using the redshift distribution derived in this work in red. We also show the Planck Collaboration et al. (2020) $T$,TE,EE+lowE results in purple. The two-dimensional histogram shows the $68\%$ and $95\%$ containment ranges and has been smoothed for visual clarity.
@@
Despite these differences in the derived redshift distributions, we do not find significant changes in the cosmological constraints from cosmic shear. In Fig. 7, we show the posterior constraints on $\Omega_{\mathrm{m}}$, the matter density, and $S_{8}=\sigma_{8}\sqrt{\Omega_{\mathrm{m}}/0.3}$, where $\sigma_{8}$ is the amplitude of matter fluctuations. The results come from a re-analysis of COSEBIs (Complete Orthogonal Sets of E/B-Integrals) in the KiDS-1000 cosmic shear data set using the CosmosSIS pipeline (Zuntz et al., 2015), as described in Asgari et al. (2021). The only noteworthy difference is that we use the nautilus Bayesian sampler (Lange, 2023) instead of MultiNest (Feroz et al., 2009) to compute the posterior, leading to more accurate but slightly weaker posterior constraints (Lemos et al., 2023). In Fig. 7, we compare the fiducial results using the redshift calibration of Hildebrandt et al. (2021) with those derived from our updated photometric redshift calibration. Note that the original cosmic shear analysis of Asgari et al. (2021) assumes that the redshift calibration may be biased in the mean redshift with Gaussian priors on the offset. Whereas the fiducial result uses priors with non-zero offsets derived from MICE mock catalogs, as described in Hildebrandt et al. (2021), the calibration offset priors are zero-centered for our updated analysis. Overall, we do not find strong differences in the derived constraints. While the derived constraint on $S_{8}$ is lower by $\sim 0.01$, the change happens mostly along the $S_{8}$-$\Omega_{\mathrm{m}}$-degeneracy direction and, thus, has little impact on the tension between weak gravitational lensing and the Planck Collaboration et al. (2020) CMB results, the so-called $S_{8}$-tension (Abdalla et al., 2022).
6.3 DESI COSMOS-XMM
@@
In this work, we have presented the first unified photometric redshift calibration of weak lensing data with DESI redshifts. We find that the single DESI data set is already competitive with respect to the entirety of previous spectroscopic redshift data sets combined. For example, the DES Y3 analysis by Myles et al. (2021) utilized spectroscopic redshifts from zCOSMOS (Lilly et al., 2009), C3R2 (Masters et al., 2019), VVDS (Le Fèvre et al., 2013), and VIPERS (Scodeggio et al., 2018) and was able to calibrate up to $47\%$ of the weak lensing source redshift distribution with spectroscopic redshifts. In contrast, in this analysis for $p_{\mathrm{f,min}}=0.1$, we are able to calibrate $69\%$ of DES with DESI redshifts alone. Subsequent iterations of the DESI COSMOS-XMM catalogs will likely increase this fraction further, making DESI an ideal data set for redshift calibration. In particular, our neural network fiber assignment probabilities allow us to easily identify bright, easy-to-observe galaxies that are currently undersampled in DESI COSMOS-XMM. At the same time, DESI’s coverage is currently poor for fainter source galaxies, as indicated by the much lower spectroscopic fractions for the HSC samples. Additionally, near-IR follow-up observations may be needed for high-redshift objects for which redshifts are hard to obtain within DESI’s wavelength coverage, $360$ nm to $980$ nm. Thus, calibrating a significant fraction of weak lensing source galaxies from the Vera C. Rubin Observatory down to faint magnitudes would require additional deep observations. In the future, we plan to supplement DESI redshifts with other spectroscopic samples such as zCOSMOS. Finally, we want to point out that the footprint of the DESI COSMOS-XMM catalog extends far beyond the original, roughly $1\,\mathrm{deg}^{2}$ COSMOS footprint and the corresponding DES, HSC, and KiDS COSMOS footprints, as shown in Fig. 1. Thus, extending photometric data to better match the DESI COSMOS-XMM spectroscopy footprint, as is expected with the Rubin deep drilling fields (Scolnic et al., 2018), will lead to a substantial reduction in cosmic variance and shape noise.
7 Conclusion
@@
  • We find excellent agreement with previously published redshift calibration results for DES Y3 (Myles et al., 2021) and HSC Y1 (Hikage et al., 2019). On the other hand, more significant differences exists with respect to the results of the fiducial KiDS-1000 analysis (Wright et al., 2020; Hildebrandt et al., 2021). We speculate that this difference may be related the use of COSMOS2020 multi-band redshifts, the peculiar photometry in the KiDS COSMOS field, and subtle statistical biases in the use of quality cuts.
@@
Our analysis method and data are also readily applicable to other weak lensing data sets such as the weak lensing catalogs from the Sloan Digital Sky Survey (SDSS; Mandelbaum et al., 2005, 2013) or the Dark Energy Camera Legacy Survey (DECaLS; Yao et al., 2023). Most importantly, we find that DESI is an excellent instrument for the photometric redshift calibration of future data sets such as the weak lensing catalog from the Vera C. Rubin Observatory. To fully achieve this potential, we also showed that larger-area deep fields matching the DESI focal plane footprint are needed.
Recent cosmological tensions such as the $S_{8}$-tension (Abdalla et al., 2022) between cosmic microwave background observations and weak lensing measurements sensitively depend on redshift calibration. For the current stage-III surveys, our analysis suggests that redshift calibration is robust enough as to not strongly affect cosmology results. In the future, DESI may play a crucial role in achieving the stringent calibration requirements of stage-IV weak lensing surveys.
@@ Acknowledgments
We thank Justin Myles and Biprateep Dey for comments on the manuscript. We are also grateful to the anonymous referee whose comments improved the clarity of the paper.
This work made use of the following software: Astropy (Astropy Collaboration et al., 2013), matplotlib (Hunter, 2007), NumPy (Van Der Walt et al., 2011), SciPy (Virtanen et al., 2020), scikit-learn (Pedregosa et al., 2011), CAMB (Lewis et al., 2000; Howlett et al., 2012), and Spyder.
AHW and HHi are supported by funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 “Cosmic Interacting Matters - From Source to Signal”. AHW is also supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz. HHi is also supported by a Heisenberg grant of the Deutsche Forschungsgemeinshaft (Hi 1495/5-1), and a European Research Council Consolidator Grant (No. 770935).

### v2

@@ VI.1 Direct Calibration
VI.2 KiDS-1000 Results
Whereas our results for DES Y3 and HSC Y1 are in good agreement with the fiducial results presented in Myles2021_MNRAS_505_4249 and Hikage2019_PASJ_71_43, we find stronger differences with the KiDS-1000 results presented in Hildebrandt2021_AA_647_124, as indicated in Fig. 6 and Table 1.
One major difference in our analysis is the use of high-quality photometric redshifts from the COSMOS2020 catalog to supplement spectroscopic data. Using COSMOS2020 redshifts is necessary since some galaxies are severely under-represented in the DESI COSMOS-XMM catalog. The effect of including COSMOS2015 multi-band redshifts in the KiDS-1000 analysis was studied in vandenBusch2022_AA_664_170 using the same methodology as in Hildebrandt2021_AA_647_124. It was found that a calibration that includes COSMOS2015 redshifts leads to higher mean redshifts even for identical SOM cells. Since we also find higher mean redshifts on average, this indicates that the use of COSMOS2020 redshifts may be one of the reasons for the different results we find here. In turn, part of this may be driven by scatter and catastrophic outliers in COSMOS2020 redshifts (Weaver2022_ApJS_258_11).
As mentioned earlier, in principle, assigning extra weights to galaxies in the calibration field to reproduce the color and magnitude distribution in the entire survey is not necessary for unbiased results. This requires that the color distribution and depth of the calibration field is representative of the full survey. However, for KiDS-1000, we find that the KiDS observations in the COSMOS field are somewhat deeper than the wider survey. Whereas Hildebrandt2021_AA_647_124 implicitly make the KiDS COSMOS catalog reproduce the color and magnitude distribution in KiDS-1000 via a SOM, we did not. This may also explain some of the differences in our results. However, it is not clear that applying extra weights, as was done in Hildebrandt2021_AA_647_124, resolves all issues associated with non-representative photometry. In the appendix, we present a comparison of the color and magnitude distribution of galaxies in the KiDS COSMOS catalog compared to the wider survey.
Finally, another reason for the different results may be related to the SOM cell quality cuts that were applied in the KiDS-1000 analysis (Hildebrandt2021_AA_647_124). Before performing the photometric redshift calibration, Hildebrandt2021_AA_647_124 defined the so-called gold selection by removing galaxies for which the mean spectroscopic redshifts of the calibration sources differ systematically from the mean photometric redshifts in that SOM cell.
@@ VI.2 KiDS-1000 Results
In the above equation, the subscripts d and w denote averages of the deep calibration field and the wide survey, respectively. Also, whereas the quality cut is applied to each SOM cell individually, the normalized median absolute deviation on the right-hand side is taken over all SOM cells. The goal of this quality cut is to remove parts of the color and magnitude space for which photometric redshifts are less reliable. Using the MICE mock catalogs, Wright2020_AA_637_100 show that this is indeed the case and reduces the bias of the SOM direct calibration method.
@@
These advantages notwithstanding, the quality cut may lead to slight statistical biases. Whether any SOM cell fulfills this quality cut is a random variable and can be very stochastic given that one typically has $\mathcal{O}(10)$ galaxies with spectroscopic redshifts in each SOM cell. A potential issue is that Hildebrandt2021_AA_647_124 use the same galaxies to define the quality cut and to perform the photometric redshift calibration. In particular, by construction, all SOM cells in the gold selection pass the quality cut, i.e., do not show significant differences between the spectroscopic and photometric redshifts. For example, the analysis of Hildebrandt2021_AA_647_124 finds that for the first tomographic bin, KiDS-1, no single SOM cell has a mean redshift above $\sim 0.6$ which implies that the reconstructed $n(z)$ drops off sharply for $z\gtrsim 0.6$. However, this is, to some extent, a statistical artifact from the fact that the same sample was used to determine the quality cuts and to perform the calibration. If one uses galaxies not used to define the quality cut, as was done in this work, it is possible that individual SOM cells may have mean redshifts above $\sim 0.6$. This will lead to a higher $n(z)$ at high redshifts, as we see in the comparison of the reconstructed $n(z)$ for KiDS-1. In summary, the implementation of the quality cuts in Hildebrandt2021_AA_647_124 may introduce subtle statistical biases which may also be partially responsible for the different results.
@@
Figure 7: Posterior constraints on $S_{8}$ and $\Omega_{\mathrm{m}}$ from the KiDS-1000 cosmic shear analysis of COSEBIs. In blue, we show the fiducial results from Asgari2021_AA_645_104, whereas we show the updated results using the redshift distribution derived in this work in red. We also show the PlanckCollaboration2020_AA_641_6 $T$,TE,EE+lowE results in purple. The two-dimensional histogram shows the $68\%$ and $99\%$ containment ranges and has been smoothed for visual clarity.
@@
Despite these differences in the derived redshift distributions, we do not find significant changes in the cosmological constraints from cosmic shear. In Fig. 7, we show the posterior constraints on $\Omega_{\mathrm{m}}$, the matter density, and $S_{8}=\sigma_{8}\sqrt{\Omega_{\mathrm{m}}/0.3}$, where $\sigma_{8}$ is the amplitude of matter fluctuations. The results come from a re-analysis of COSEBIs (Complete Orthogonal Sets of E/B-Integrals) in the KiDS-1000 cosmic shear data set using the CosmosSIS pipeline (Zuntz2015_AC_12_45), as described in Asgari2021_AA_645_104. The only noteworthy difference is that we use the nautilus Bayesian sampler (Lange2023_MNRAS_525_3181) instead of MultiNest (Feroz2009_MNRAS_398_1601) to compute the posterior, leading to more accurate but slightly weaker posterior constraints (Lemos2023_MNRAS_521_1184). In Fig. 7, we compare the fiducial results using the redshift calibration of Hildebrandt2021_AA_647_124 with those derived from our updated photometric redshift calibration. Note that the original cosmic shear analysis of Asgari2021_AA_645_104 assumes that the redshift calibration may be biased in the mean redshift with Gaussian priors on the offset. Whereas the fiducial result uses priors with non-zero offsets derived from MICE mock catalogs, as described in Hildebrandt2021_AA_647_124, the calibration offset priors are zero-centered for our updated analysis. Overall, we do not find strong differences in the derived constraints. While the derived constraint on $S_{8}$ is lower by $\sim 0.01$, the change happens mostly along the $S_{8}$-$\Omega_{\mathrm{m}}$-degeneracy direction and, thus, has little impact on the tension between weak gravitational lensing and the PlanckCollaboration2020_AA_641_6 CMB results, the so-called $S_{8}$-tension (Abdalla2022_JHEAp_34_49).
VI.3 DESI COSMOS-XMM
@@
In this work, we have presented the first unified photometric redshift calibration of weak lensing data with DESI redshifts. We find that the single DESI data set is already competitive with respect to the entirety of previous spectroscopic redshift data sets combined. For example, the DES Y3 analysis by Myles2021_MNRAS_505_4249 utilized spectroscopic redshifts from zCOSMOS (Lilly2009_ApJS_184_218), C3R2 (Masters2019_ApJ_877_81), VVDS (LeFevre2013_AA_559_14), and VIPERS (Scodeggio2018_AA_609_84) and was able to calibrate up to $47\%$ of the weak lensing source redshift distribution with spectroscopic redshifts. In contrast, in this analysis for $p_{\mathrm{f,min}}=0.1$, we are able to calibrate $69\%$ of DES with DESI redshifts alone. Subsequent iterations of the DESI COSMOS-XMM catalogs will likely increase this fraction further, making DESI an ideal data set for redshift calibration. In particular, our neural network fiber assignment probabilities allow us to easily identify bright, easy-to-observe galaxies that are currently undersampled in DESI COSMOS-XMM. At the same time, DESI’s coverage is currently poor for fainter source galaxies, as indicated by the much lower spectroscopic fractions for the HSC samples. Additionally, near-IR follow-up observations may be needed for high-redshift objects for which redshifts are hard to obtain within DESI’s wavelength coverage, $360$ nm to $980$ nm. Thus, calibrating a significant fraction of weak lensing source galaxies from the Vera C. Rubin Observatory down to faint magnitudes would require additional deep observations. In the future, we plan to supplement DESI redshifts with other spectroscopic samples such as zCOSMOS. Finally, we want to point out that the footprint of the DESI COSMOS-XMM catalog extends far beyond the original, roughly $1\,\mathrm{deg}^{2}$ COSMOS footprint and the corresponding DES, HSC, and KiDS COSMOS footprints, as shown in Fig. 1. Thus, extending photometric data to better match the DESI COSMOS-XMM spectroscopy footprint, as is expected with the Rubin deep drilling fields (Scolnic2018_arXiv_1812_0516), will lead to a substantial reduction in cosmic variance and shape noise.
VII Conclusion
@@
  • We find excellent agreement with previously published redshift calibration results for DES Y3 (Myles2021_MNRAS_505_4249) and HSC Y1 (Hikage2019_PASJ_71_43). On the other hand, more significant differences exists with respect to the results of the fiducial KiDS-1000 analysis (Wright2020_AA_637_100; Hildebrandt2021_AA_647_124). We speculate that this difference may be related the use of COSMOS2020 multi-band redshifts, the peculiar photometry in the KiDS COSMOS field, and subtle statistical biases in the use of quality cuts.
@@
Our analysis method and data are also readily applicable to other weak lensing data sets such as the weak lensing catalogs from the Sloan Digital Sky Survey (SDSS; Mandelbaum2005_MNRAS_361_1287; Mandelbaum2013_MNRAS_432_1544) or the Dark Energy Camera Legacy Survey (DECaLS; Yao2023_MNRAS_524_6071). Most importantly, we find that DESI is an excellent instrument for the photometric redshift calibration of future data sets such as the weak lensing catalog from the Vera C. Rubin Observatory. To fully achieve this potential, we also showed that larger-area deep fields matching the DESI focal plane footprint are needed.
Recent cosmological tensions such as the $S_{8}$-tension (Abdalla2022_JHEAp_34_49) between cosmic microwave background observations and weak lensing measurements sensitively depend on redshift calibration. For the current stage-III surveys, our analysis suggests that redshift calibration is robust enough as to not strongly affect cosmology results. In the future, DESI may play a crucial role in achieving the stringent calibration requirements of stage-IV weak lensing surveys.
@@ Acknowledgments
We thank Justin Myles and Biprateep Dey for comments on the manuscript. We are also grateful to the anonymous referee whose detailed and thorough report improved the clarity of the paper.
This work made use of the following software: Astropy (AstropyCollaboration2013_AA_558_33), matplotlib (Hunter2007_CSE_9_90), NumPy (VanDerWalt2011_CSE_13_22), SciPy (Virtanen2020_NatMe_17_261), scikit-learn (Pedregosa2011_JMLR_12_2825), CAMB (Lewis2000_ApJ_538_473; Howlett2012_JCAP_04_027), and Spyder.
AHW and HHi are supported by funding from the German Science Foundation DFG, via the Collaborative Research Center SFB1491 “Cosmic Interacting Matters - From Source to Signal”. AHW is also supported by the Deutsches Zentrum für Luft- und Raumfahrt (DLR), made possible by the Bundesministerium für Wirtschaft und Klimaschutz. HHi is also supported by a Heisenberg grant of the Deutsche Forschungsgemeinshaft (Hi 1495/5-1), and a European Research Council Consolidator Grant (No. 770935). Work by AF has been supported by a CAS Summer Undergraduate Fellowship from American University’s College of Arts & Sciences (CAS) and the NASA DC Space Grant Consortium (DCSGC).
@@
Our analysis method inherently assumes that the properties of galaxies, in the deep fields, including their measured photometry, are representative of the wider weak lensing survey. However, this may not be the case for KiDS-1000. One reason for this is that the KiDS pointings have a range of depths in the ugriugri bands and the COSMOS pointing may not perfectly representative of the mean. For example, while the $r$-band depth of COSMOS pointing is within the range of KiDS-1000 pointings, it is deeper than the average. On the other hand, the $ZYJHK_{s}$ bands were actually observed at higher depth but artificially downgraded by adding gaussian noise. Ultimately, for KiDS-1000, we find statistically significant differences in the color and magnitude distribution of KiDS observations in COSMOS deep field compared to KiDS-1000.
Figure 8: The effective number of galaxies in each SOM cell for the full KiDS SOM-gold sample (left), the transfer catalog used in this work (middle), and the transfer catalog used in Hildebrandt2021_AA_647_124. The values are normalized such that the average number over all SOM cells is 1 in all cases. Neither transfer catalog faithfully reproduces all features of the KiDS SOM-gold catalog.
Fig. 8 shows the color distribution of galaxies in the KiDS SOM, weighted by the lensing weight $w_{\mathrm{l}}$. We compare the results from KiDS-1000 against our COSMOS catalog using the weights derived from neural networks and the fiducial KiDS COSMOS catalog using weights derived from nearest-neighbor matching (Hildebrandt2021_AA_647_124). While the KiDS COSMOS catalog reproduce many salient features in the SOM map, clear differences also emerge, regardless of the assumed weight in KiDS COSMOS. For example, one can see that the distribution of galaxies in the KiDS COSMOS catalog reaches to fainter $r$-band magnitudes than the KiDS-1000 gold catalog.
Figure 9: Weighted distribution of $r$ magnitudes and rZr-Z colors of galaxies in the KiDS SOM-gold sample (grey) compared to the KiDS COSMOS catalog used in this work (red) and the fiducial KiDS COSMOS catalog (blue). We show $99\%$, $95\%$, and $68\%$ ranges. The densities have been smoothed with a Gaussian kernel with a width of $0.2$ mag for visual clarity.
This is further investigated in Fig. 9, where we show the corresponding $r$ vs. rZr-Z color and magnitude distributions. The KiDS COSMOS catalog using nearest-neighbor weights clearly shows substantial differences to KiDS-1000. The main reason is that these weights do not account for the probability $w_{\mathrm{t}}(s)$ that a galaxy makes the KiDS-1000 weak lensing cuts. However, even for the neural network weights that account for $w_{\mathrm{t}}(s)$, the KiDS COSMOS catalog extends to deeper magnitudes and shows redder colors than KiDS-1000. We observe similar color and magnitude trends in the raw catalogs, i.e., before applying lensing weights, indicating that this is not due to our approximate method for creating KiDS-1000-like lensing weights. All of this indicates that the KiDS COSMOS catalog is not perfectly representative of the KiDS-1000 data set and may lead to biases in the estimated redshift distribution.